# GraphRAFT: Retrieval Augmented Fine-Tuning for Knowledge Graphs on Graph Databases

### v1

@@ Abstract
Large language models have shown remarkable language processing and reasoning ability but are prone to hallucinate when asked about private data. Retrieval-augmented generation (RAG) retrieves relevant data that fit into an LLM’s context window and prompts the LLM for an answer. GraphRAG extends this approach to structured Knowledge Graphs (KGs) and questions regarding entities multiple hops away. The majority of recent GraphRAG methods either overlook the retrieval step or have ad hoc retrieval processes that are abstract or inefficient. This prevents them from being adopted when the KGs are stored in graph databases supporting graph query languages. In this work, we present GraphRAFT, a retrieve-and-reason framework that finetunes LLMs to generate provably correct Cypher queries to retrieve high-quality subgraph contexts and produce accurate answers. Our method is the first such solution that can be taken off-the-shelf and used on KGs stored in native graph DBs. Benchmarks suggest that our method is sample-efficient and scales with the availability of training data. Our method achieves significantly better results than all state-of-the-art models across all four standard metrics on two challenging Q&As on large text-attributed KGs.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable ability to reason over natural languages. However, when prompted with questions regarding new or private knowledge, LLMs are prone to hallucinations[1]. A popular approach to remediate this issue is Retrieval-Augmented Generation (RAG) [2]. Given a natural language question and a set of text documents, RAG retrieves a small set of relevant documents and feeds them into an LLM.
@@
In addition to unstructured data such as text documents, many of the real-world structured data come in the form of Knowledge Graphs (KGs). KGs can be used to model a variety of domains such as the Web, social networks, and financial systems. Knowledge Base Question Answering (KBQA)[3] studies the methods used to answer questions about KGs. Most of the work studies multi-hop questions on the Web (open-domain)[4, 5, 6].
@@
Historically, there are two categories of approaches for KBQA. One approach is based on semantic parsing[7, 8, 9, 10, 11, 12, 13], which involves converting natural language questions into logical forms that are often SPARQL[14]. Another approach is based on information retrieval[15, 16, 17, 18], which involves embedding and retrieving data (mostly triplets) from KGs and reranking the extracted data.
@@
There is a growing demand to leverage the language understanding and reasoning power of LLMs to improve open-domain KBQA. At the same time, GraphRAG extends the RAG setup to private KGs. The methodologies began to merge, despite the different application areas. The vast majority of recent literature follows the general framework of 1) identifying some entities in the given question, 2) retrieving a subgraph and 3) prompting an LLM for reasoning.
@@
[19, 20] retrieves and reranks the top nodes (or k-hop ego-subgraphs) and prompts the LLM. [21] prompts an LLM to generate relation path templates and retrieves concrete paths using in-memory breadth-first search and fintunes the LLM. [22, 23, 24, 25, 26] identify starting entities and iteratively perform retrieval step-by-step by prompting the LLM what is the next step to take. [27] prunes a retrieved subgraph with parameterised graph algorithms and prompt an LLM with a textualised version of the local subgraph. [28] extends LLM chain-of-thought (CoT) by iteratively retrieving from the KG. [29] enhances the quality of the QA training data.
@@
Although significant progress has been made, there are several limitations of existing methods. Many of the real-world KGs are stored in native graph DBs. Graph DBs come with query engines that optimise user queries, such as Cypher. However, none of the existing methods leverage such graph DBs at all. Any method that explicitly requires step-by-step retrieval (due to iteratively prompting the LLMs) prevents direct multi-hop querying on DBs. This implies that each step of the retrieval process needs to be fully materialised and the fixed sequential order of traversal implementation is frequently suboptimal. Benchmarking are often performed with in-memory graphs represented as tensors with adhoc implementations of subgraph retrieval.
@@
On the other hand, many methods that involve multi-hop retrieval as a single step remain abstract. They usually stop at the point of specifying some path patterns. Converting the abstract path definitions into queries remains a separate challenging task, reardless of specific query languages[30, 31, 32, 33]. Another stream of work explicitly requires the questions to be of some logic fragment and converts them into SPARQL. There is also no general way of guaranteeing the generated queries are syntactically and semantically correct with respect to specific KGs.
@@
In this paper, we propose a simple, modular and reliable approach for question answering on private KGs (GraphRAG). Specifically, the contributions of our papers are:
@@
  • We propose a new method of finetuning an LLM to generate optimal Cyphers that only requires textual QAs as training data but not ground-truth queries.
@@
In addition, benchmarks on text-rich KGs suitable for GraphRAG show our method achieves significantly beyond SOTA results across all 4 metrics on STaRK-prime and STaRK-mag, two large text-attributed KGs with Q&As for GraphRAG. For example, our method achieves 63.71% Hit@1 and 68.99% Mean Reciprocal Rank (MRR), which are 22.81% and 17.79% better than the best previous baseline respectively on STaRK-prime. Our method is sample-efficient, for example achieving beyond SOTA results on all metrics on STaRK-prime when trained with only 10% of the data, and scales with more training data.
@@ 2 Related Work
RAG. Recent work extends RAG[2] to broader settings. RAFT[34] studies domain-specific retrieval-augmented fine-tuning on by sampling relevant and irrelavant documents. GraphRAG[35] generalises RAG to global text summarisation tasks in multiple documents. The term GraphRAG has also been widely adopted to mean RAG on (knowledge) graphs, which is the problem setup we study in this paper. [36] applies RAFT for GraphRAG on document retrieval and summarisation. There are many methodological and implementation improvements[37] and industrial applications of GraphRAG[38].
Message Passing Graph Neural Networks and Graph Transformers. Message-passing Graph Neural Networks (GNNs) iteratively aggregates and updates embeddings on nodes and edges using the underlying graph as the computation graph, hence incorporating strong inductive bias. GNNs have been used within the framework of GraphRAG to improve retrieval[39] or to enhance the graph processing power of LLMs[40, 41]. Alternative model architectures such as Graph Transformers, which employ graph-specific tokenization or positional embedding techniques, have been developed[42, 43]. There are also work[44, 45, 46] that leverage LLMs to improve classical graph problems such as node classification and link prediction on text-attributed graphs.
@@ 3.1 Graph database and graph query language
A Graph DB stores LPGs efficiently on disks. It also comes with a query engine that allows one to query the graph. A widely used query language is Cypher[47][48] (and a variant openCypher[49]), which are implementations of the recent GQL[50][51] standard. We will refer to the two interchangeably throughout this paper. The key ingredient of Cypher is graph pattern matching. A user can query the graph by matching on patterns (e.g paths of a certain length) and predicates (e.g filtering on node and edge properties). An example Cypher query is provided in Figure 1. The execution of a query is carried out by the query engine which heavily optimises the order of executing low-level operators.
An alternative graph data model is Resource Description Framework (RDF)[52]. It was originally created to model the Web. A graph is modeled as a collection of triples, commonly referred to as subject-predicate-object, where each element can be identified with a URI (mimicing the Web). Query languages on top of RDFs include SPARQL[14]. RDFs also support ontology languages such as Web Ontology Language (OWL)[53] to model and derive facts through formal logic. Much of the KBQA literature has taken inspiration from RDFs by defining a graph as a collection of triples and reasoning over it in a formal style. Widely used benchmarks such as WebQSP[4] are exactly questions about the Web that are answerable by SPARQL.
Many open-source and commercial graph DBs support Cypher over LPGs, such as Neo4j[54], ArangoDB[55], TigerGraph[56]. There are also popular RDF stores such as Neptune[57], which also supports Cypher over RDFs.
@@
MATCH (s {name: source_name})-[r1]-(var)-[r2]-t WHERE tgt <> s
RETURN labels(s)[1] AS label1, s.name AS name1, type(r1) AS type1,
labels(var)[1] AS label2, type(r2) AS type2, labels(t)t AS label3,
@@
Figure 1: An example Cypher query. It takes as user input list variables source_names. It iterates through them and find all two-hop neighbours of each source_name node, requiring the second-hop node to be distinct to the source. The query returns aggregate information of the subgraph such as labels and types of nodes and edges, and arithmetic over how many second-hop nodes have ids that are in the user-defined node id list.
@@ 3.2 LLM
Modern LLMs are usually auto-regressive decoder-only Transformers as backbones that are trained on the Web. An LLM $f_{\theta}$ has a fixed vocabulary set $\mathcal{V}$. Given a sequence of characters $c_{0},\cdots,c_{n}$, a tokenizer converts it to a sequence of tokens $t_{0},\cdots,t_{k}$ where $t_{i}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension. The tokenizer often compresses multiple characters into one token and splits a word into multiple tokens. Given $t_{0},\cdots,t_{k}$, suppose $f_{\theta}$ has generated tokens $t_{k+1},\cdots,t_{k+n}$, it computes the logits for the $k+n+1^{\text{th}}$ token as $l_{k+n+1}=f_{\theta}(t_{0};t_{k+n})$ where $l_{k+n+1}\in\mathbb{R}^{|\mathcal{V}|}$. The probability of generating any token $x$ is obtained by applying softmax to the logits, $p(t_{k+1}=x|t_{0};t_{k+n+1})=exp(l_{k+n+1}^{x})/\sum_{y\in\mathcal{V}}{exp(l_{k+n+1}^{y})}$. Greedy decoding picks the token with the highest probability at each step. Beam search with width $m$ keeps $m$ sequence of tokens with the highest product of probabilities so far. The generative process normally terminates when some end-of-sequence $<eos>$ token is decoded. Pretrained LLMs can be finetuned efficiently using techniques such as LoRA[58] optimising the product of conditional probabilities of next-token prediction.
@@ 4 Approach
Given a set of training QAs $\{Q_{i},A_{i}\}$ where $A_{i}\subset V$, we want our model to produce good answers $A_{j}$ for unseen questions ${Q_{j}}$ according to a variety of metrics. Our approach consists of several steps. First, we create a set of training question-Cypher pairs $\{Q_{i},C_{i}\}$ to finetune a LLM (Section 4.1). At inference time, we deploy a simple method of constrained decoding that is grounded to $\mathcal{G}$ to guarantee syntactically and semantically optimal Cypher (Section 4.2). The optimal Cypher is executed to retrieve a text-attributed subgraph for each question and a second LLM is finetuned to jointly reason over text and subgraph topology to select the final answers (Section 4.3).
@@ 4.1 Synthesize ground-truth cypher
This enables more accurate identification of entities (that correspond to nodes in the graph) than performing native vector similarity between all nodes in the graph and an embedded vector representing the question.
We then construct several Cypher templates that pattern matches multi-hop graphs around the identified entities $v_{1},\cdots,v_{k}$ with filters on node and relationship types. For each Cypher query $C_{j}$ we execute the query and compute the hits in answers ${A_{i}}$, as well as the number of total nodes. The calculation (as an aggregation step) is performed as part of the query itself which means the nodes and edges and their properties do not need to be materialised and retrieved and hence saves memory and I/O workload. An example such query is shown in Figure 1.
We then rank the queries according to their hits and number of results and retains the top query $C_{i}$. This gives us a set of $\{Q_{i},C_{i}\}$ between questions and best Cypher query that contains the answers $\{A_{i}\}$. The best Cypher can be decided by reranking and for example filtering by recall and precision. We finetune an LLM $L_{1}$ with the given set of training and validation data. The workflow is illustrated in Figure 2.
@@ 4.2 Grounded constrained decoding
Our $L_{1}$ has learned to generate graph pattern matching Cypher queries. However, there is still no guarantee that the generated Cypher is executable at inference time. The major bottlenecks are 1) syntactical correctness and 2) semantic correctness (i.e any type filters must correspond to existing node labels or edge types of $\mathcal{G}$). A common approach of generate-and-validate is both costly and slow and requires a sophisticated correction scheme post-validation to guarantee eventual correctness.
We use a new method of next-token constrained decoding at the level of logits processor at inference time. The constraints are applied at the token-level instead of word (or query keyword) level since the tokenizers used by the majorioty of modern LLMs do not have a simple mapping between words and query keywords to the tokens.
Given a question at inference time, we first create a set of possible m-hop queries involving identified entities. This step is efficient since the schema of the graph is available. Let there be $\mathcal{Q}=\{\mathcal{Q}^{1},\cdots,\mathcal{Q}^{M}\}$ valid queries, each has tokenization $\mathcal{Q}^{k}=(n_{0},\cdots,n_{Q_{k}})\in\mathcal{V}^{Q_{k}}$ of variable length. When $L_{1}$ has generated $i$ tokens $[t_{0},\cdots t_{i}]$ and is generating a vector of logits $(l_{0},\cdots,l_{\mathcal{V}})$, our logits processor masks all invalid token with the value $\infty$- ∞ by comparing with the $i+1^{\text{th}}$ token of all possible tokenized queries that match the initial $i$ tokens. The masking is defined as Equation 2 where $M_{i+1}$ is the set of valid tokens at $i+1^{\text{th}}$ position grounded in $\mathcal{G}$ via $\mathcal{Q}$.
~$\tilde{\ell}_{i+1}^{(x)}=\begin{cases}\ell_{i+1}^{(x)},&\text{if }x\in M_{i+1}\\ -\infty,&\text{if }x\notin M_{i+1}\end{cases}\text{where }M_{i+1}=\{t\in\mathcal{V}|\exists\mathcal{Q}^{k}\in\mathcal{Q}\text{ , }\mathcal{Q}^{k}_{0:i+1}=(t_{0},\cdots,t_{k},t)\}$ (2)
The tokenizer for $L_{1}$ then applies decoding after softmax. By construction, our decoded query is executable. This constraint decoding is non-invasive for both greedy decoding and beam search with sufficiently large beams. Unlikely existing constrained generation methods[59, 60], our approach does not require any formal grammar to be defined and is context-aware with respect to $\mathcal{G}$. An example of grounded constrained decoding is illustrated in Figure 3. Since our method is applied entirely on the logits and before softmax, various ways of decoding such as greedy and beam search can be used independently.

### v2

@@ Abstract
Large language models have shown remarkable language processing and reasoning ability but are prone to hallucinate when asked about private data. Retrieval-augmented generation (RAG) retrieves relevant data that fit into an LLM’s context window and prompts the LLM for an answer. GraphRAG extends this approach to structured Knowledge Graphs (KGs) and questions regarding entities multiple hops away. The majority of recent GraphRAG methods either overlook the retrieval step or have ad hoc retrieval processes that are abstract or inefficient. This prevents them from being adopted when the KGs are stored in graph databases supporting graph query languages. In this work, we present GraphRAFT, a retrieve-and-reason framework that finetunes LLMs to generate provably correct Cypher queries to retrieve high-quality subgraph contexts and produce accurate answers. Our method is the first such solution that can be taken off-the-shelf and used on KGs stored in native graph DBs. Benchmarks suggest that our method is sample-efficient and scales with the availability of training data. Our method achieves significantly better results than all state-of-the-art methods across all four standard metrics on two set of challenging Q&As on large text-attributed KGs.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable ability to reason over natural languages. However, when prompted with questions regarding new or private knowledge, LLMs are prone to hallucinations[huang2025survey]. A popular approach to remediate this issue is Retrieval-Augmented Generation (RAG) [lewis2020retrieval]. Given a natural language question and a set of text documents, RAG retrieves a small set of relevant documents and feeds them into an LLM.
@@
In addition to unstructured data such as text documents, many of the real-world structured data come in the form of Knowledge Graphs (KGs). KGs can be used to model a variety of domains such as the Web, social networks, and financial systems. Knowledge Base Question Answering (KBQA)[lan2021survey] studies the methods used to answer questions about KGs. Most of the work studies multi-hop questions on the Web (open-domain)[Yih2016TheVO, freebase, talmor-berant-2018-web].
@@
Historically, there are two categories of approaches for KBQA. One approach is based on semantic parsing[yih-etal-2015-semantic, luo-etal-2018-knowledge, 10.1007/978-3-319-93417-4_46, 10.1145/2806416.2806472, chen-etal-2019-uhop, 10.1145/3357384.3358033, lan-jiang-2020-query], which involves converting natural language questions into logical forms that are often SPARQL[sparql]. Another approach is based on information retrieval[sun2019pullnetopendomainquestion, sun-etal-2018-open, xiong-etal-2019-improving, he-etal-2017-generating], which involves embedding and retrieving triplets from KGs and reranking the extracted data.
@@
There is a growing demand to leverage the language understanding and reasoning power of LLMs to improve open-domain KBQA. At the same time, GraphRAG extends the RAG setup to private KGs. The methodologies began to merge, despite the different application areas. A unifying architecture follows the general framework of 1) identifying some entities in the given question, 2) retrieving a subgraph and 3) prompting an LLM for reasoning.
@@
[baek2023knowledgeaugmentedlanguagemodelprompting, hu2024graggraphretrievalaugmentedgeneration] retrieves and reranks the top nodes (or k-hop ego-subgraphs) and prompts the LLM. [luo2024reasoning] prompts an LLM to generate relation path templates and retrieves concrete paths using in-memory breadth-first search and fintunes the LLM. [sun2024thinkongraph, ma2024think, sui2024fidelisfaithfulreasoninglarge, wang2025reasoning, xia2024knowledgeawarequeryexpansionlarge] identify starting entities and iteratively perform the retrieval step-by-step by prompting the LLM what is the next step to take. [he2024gretriever] prunes a given subgraph with parameterised graph algorithms and prompts an LLM with a textualised version of the local subgraph. [wang2023knowledgedrivencotexploringfaithful] extends LLM chain-of-thought (CoT) by iteratively retrieving from the KG. [chen2025knowledge] enhances the quality of the QA training data.
@@
Although significant progress has been made, there are several limitations of existing methods. Many of the real-world KGs are stored in native graph DBs. Graph DBs come with query engines that optimise user queries, such as Cypher. However, none of the existing methods leverage such graph DBs at all. Any method that explicitly requires step-by-step retrieval (due to iteratively prompting the LLMs) prevents direct multi-hop querying on DBs. This implies that each step of the retrieval process needs to be fully materialised and the fixed sequential order of traversal implementation is frequently suboptimal. Benchmarking is often performed with in-memory graphs represented as tensors with adhoc implementations of subgraph retrieval.
@@
On the other hand, many methods that involve multi-hop retrieval as a single step remain abstract. They usually stop at the point of specifying some path patterns. Converting the abstract path definitions into queries remains a separate challenging task, irrespective of specific query languages[lei2025spider20evaluatinglanguage, texttosparql, brei2024leveragingsmalllanguagemodels, ozsoy2024text2cypherbridgingnaturallanguage]. Another stream of work explicitly requires the questions to be of some logic fragment and converts them into SPARQL. There is also no general way of guaranteeing that the generated queries are syntactically and semantically correct with respect to specific KGs.
@@
In this paper, we propose a simple, modular, and reliable approach for question answering on private KGs (GraphRAG). Specifically, the contributions of our papers are:
@@
  • We propose a new method of fine-tuning an LLM to generate optimal Cypher queries that only requires textual QAs as training data but not user-crafted ground-truth queries.
@@
In addition, benchmarks show that our method achieves significantly beyond SOTA results across all 4 metrics on STaRK-prime and STaRK-mag, two large text-attributed KGs with Q&As for GraphRAG. For example, our method achieves 63.71% Hit@1 and 68.99% Mean Reciprocal Rank (MRR), which is an improvement of 22.81 and 17.79 percentage points, respectively, over the best previous baselines on STaRK-prime. Our method is sample-efficient, achieving beyond SOTA results on all metrics on STaRK-prime when trained with only 10% of the data, and scales with more training data.
@@ 2 Related Work
RAG. Recent work extends RAG[lewis2020retrieval] to broader settings. RAFT[raft2024] studies domain-specific retrieval-augmented fine-tuning on by sampling relevant and irrelavant documents. GraphRAG[edge2024local] generalises RAG to global text summarisation tasks in multiple documents. The term GraphRAG has also been widely adopted to mean RAG on (knowledge) graphs, which is the problem setup we study in this paper. [graft] applies RAFT for GraphRAG on document retrieval and summarisation. There are many methodological and implementation improvements[li2024simple] and industrial applications of GraphRAG[Xu_2024].
Message Passing Graph Neural Networks and Graph Transformers. Message-passing Graph Neural Networks (GNNs) iteratively aggregates and updates embeddings on nodes and edges using the underlying graph as the computation graph, hence incorporating strong inductive bias. GNNs have been used within the framework of GraphRAG to improve retrieval[mavromatis2024gnn] or to enhance the graph processing power of LLMs[zhang2022greaselm, yasunaga2021qa]. Alternative model architectures such as Graph Transformers, which employ graph-specific tokenization or positional embedding techniques, have been developed[ying2021transformersreallyperformbad, tokengt]. There are also work[zhao2023learning, he2024harnessing, jiang2024ragraph] that leverage LLMs to improve classical graph problems such as node classification and link prediction on text-attributed graphs.
@@ 3.1 Graph database and graph query language
A Graph DB stores LPGs efficiently on disks. It also comes with a query engine that allows one to query the graph. A widely used query language is Cypher[cypher][cyphersemantics] (and a variant openCypher[opencypher]), which are implementations of the recent GQL[gql][gqlpaper] standard. We will refer to the two interchangeably throughout this paper. The key ingredient of Cypher is graph pattern matching. A user can query the graph by matching on patterns (e.g paths of a certain length) and predicates (e.g filtering on node and edge properties). An example Cypher query is provided in Figure 1. The execution of a query is carried out by the query engine which heavily optimises the order of executing low-level operators.
An alternative graph data model is Resource Description Framework (RDF)[rdf]. It was originally created to model the Web. A graph is modeled as a collection of triples, commonly referred to as subject-predicate-object, where each element can be identified with a URI (mimicing the Web). Query languages on top of RDFs include SPARQL[sparql]. RDFs also support ontology languages such as Web Ontology Language (OWL)[owl] to model and derive facts through formal logic. Much of the KBQA literature has taken inspiration from RDFs by defining a graph as a collection of triples and reasoning over it in a formal style. Widely used benchmarks such as WebQSP[Yih2016TheVO] are exactly questions about the Web that are answerable by SPARQL.
Many open-source and commercial graph DBs support Cypher over LPGs, such as Neo4j[neo4j], ArangoDB[arangoDB], TigerGraph[tigergraph]. There are also popular RDF stores such as Neptune[neptune], which also supports Cypher over RDFs.
@@
MATCH (src {name: source_name})-[r1]-(var)-[r2]-(tgt) WHERE src <> tgt
RETURN labels(src)[1] AS label1, src.name AS name1, type(r1) AS type1,
labels(var)[1] AS label2, type(r2) AS type2, labels(tgt)[1] AS label3,
@@
Figure 1: An example Cypher query. It takes as user input a list variable source_names and another list target_ids. It iterates through them and find all two-hop neighbours of each source_name node, requiring the second-hop node to be distinct to the source. The query returns aggregate information of the subgraph such as labels and types of nodes and edges, and arithmetic over how many second-hop nodes have ids that are in the user-defined node id list.
@@ 3.2 LLM
Modern LLMs are usually auto-regressive decoder-only Transformers as backbones that are trained on the Web. An LLM $f_{\theta}$ has a fixed vocabulary set $\mathcal{V}$. Given a sequence of characters $c_{0},\cdots,c_{n}$, a tokenizer converts it to a sequence of tokens $t_{0},\cdots,t_{k}$ where $t_{i}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension. The tokenizer often compresses multiple characters into one token and splits a word into multiple tokens. Given $t_{0},\cdots,t_{k}$, suppose $f_{\theta}$ has generated tokens $t_{k+1},\cdots,t_{k+n}$, it computes the logits for the $k+n+1^{\text{th}}$ token as $l_{k+n+1}=f_{\theta}(t_{0};t_{k+n})$ where $l_{k+n+1}\in\mathbb{R}^{|\mathcal{V}|}$. The probability of generating any token $x$ is obtained by applying softmax to the logits, $p(t_{k+n+1}=x|t_{0};t_{k+n})=exp(l_{k+n+1}^{x})/\sum_{y\in\mathcal{V}}{exp(l_{k+n+1}^{y})}$. Greedy decoding picks the token with the highest probability at each step. Beam search with width $m$ keeps $m$ sequences of tokens with the highest product of probabilities so far. The generative process normally terminates when some end-of-sequence <eos><eos> token is decoded. Pretrained LLMs can be finetuned efficiently using techniques such as LoRA[hu2022lora], optimising the product of conditional probabilities of next-token prediction.
@@ 4 Approach
Given a set of training QAs $\{Q_{i},A_{i}\}$ where $A_{i}\subset V$, we want our model to produce good answers $A_{j}$ for unseen questions $Q_{j}$} according to a variety of metrics. Our approach consists of several steps. First, we create a set of training question-Cypher pairs $\{Q_{i},C_{i}\}$ to finetune a LLM (Section 4.1). At inference time, we deploy a simple method of constrained decoding that is grounded to $\mathcal{G}$ to guarantee syntactically and semantically optimal Cypher queries (Section 4.2). The optimal Cypher is executed to retrieve a text-attributed subgraph for each question and a second LLM is finetuned to jointly reason over text and subgraph topology to select the final answers (Section 4.3).
@@ 4.1 Synthesize ground-truth cypher
This enables more accurate identification of entities than performing native vector similarity between all nodes in the graph and an embedded vector representing the question.
We then construct several Cypher templates that match multi-hop subgraphs around the identified entities $v_{1},\cdots,v_{k}$ with filters on node and relationship types. For each Cypher query $C_{j}$ we execute the query and compute the hits in the answers $A_{i}$}, as well as the number of total nodes. The calculation (as an aggregation step) is performed as part of the query itself which means the nodes and edges and their properties do not need to be materialised and retrieved and hence saves memory and I/O workload. An example such query is shown in Figure 1.
We then rank the queries according to their hits and number of results and retain the top query $C_{i}$. This gives us a set of $\{Q_{i},C_{i}\}$ between questions and best Cypher queries that contain the answers $\{A_{i}\}$. The best Cypher can be decided by reranking and for example filtering by recall and precision. We finetune an LLM $L_{1}$ with the given set of training data. The workflow is illustrated in Figure 2.
@@ 4.2 Grounded constrained decoding
Our $L_{1}$ has learned to generate graph pattern matching Cypher queries. However, there is still no guarantee that the generated Cypher is executable at inference time. The major bottlenecks are 1) syntactical correctness and 2) semantic correctness (i.e any type filters must correspond to existing node labels or edge types of $\mathcal{G}$). A common approach of generate-and-validate is both costly and slow and requires a sophisticated post-validation correction scheme to guarantee eventual correctness.
We use a new method of next-token constrained decoding at the level of logits processor at inference time. The constraints are applied at the token-level instead of word (or query keyword) level since the tokenizers used by the majorioty of modern LLMs do not have a one-to-one mapping from words and query keywords to the tokens.
Given a question at inference time, we first create a set of possible m-hop queries involving identified entities. This step is efficient since the schema of the graph is available. Let there be $\mathcal{Q}=\{\mathcal{Q}^{1},\cdots,\mathcal{Q}^{M}\}$ valid queries, each has tokenization $\mathcal{Q}^{k}=(n_{0},\cdots,n_{Q_{k}})\in\mathcal{V}^{Q_{k}}$ of variable length. When $L_{1}$ has generated $i$ tokens $(t_{0},\cdots,t_{i})$ and is generating a vector of logits $(l_{0},\cdots,l_{\mathcal{V}})$, our logits processor masks all invalid token with the value -$\infty$ by comparing with the $i+1^{\text{th}}$ token of all possible tokenized queries that match the initial $i$ tokens. The masking is defined as Equation 2 where $M_{i+1}$ is the set of valid tokens at $i+1^{\text{th}}$ position grounded in $\mathcal{G}$ via $\mathcal{Q}$.
$\tilde{\ell}_{i+1}^{(x)}=\begin{cases}\ell_{i+1}^{(x)},&\text{if }x\in M_{i+1}\\ -\infty,&\text{if }x\notin M_{i+1}\end{cases}\text{where }M_{i+1}=\{t\in\mathcal{V}|\exists\mathcal{Q}^{k}\in\mathcal{Q}\text{ , }\mathcal{Q}^{k}_{0:i+1}=(t_{0},\cdots,t_{k},t)\}$ (2)
The tokenizer for $L_{1}$ then applies decoding after softmax. By construction, our decoded query is executable. This constraint decoding is faithful and non-invasive for both greedy decoding and beam search with sufficiently large beams, as stated in Lemma 1. The formal statement and proof are provided in Appendix A. Unlike existing constrained generation methods[beurerkellner2024guidingllmsrightway, geng2024grammarconstraineddecodingstructurednlp], our approach does not require any formal grammar to be defined and is context-aware with respect to $\mathcal{G}$. An example of grounded constrained decoding is illustrated in Figure 3. Since our method is applied entirely on the logits before softmax, different final decoding such as greedy and beam search can be applied.
Lemma 1.
(Informal) If a query is invalid, it will not be generated. When using beam width = 1, constrained decoding acts as greedy decoding among valid queries. When beam width = M, exactly all valid queries are generated.